Let R be a reflexive relation on a finite set A, and let MR be the bit matrix representing R. Specify the value of the entries on the main diagonal.
1
Expert's answer
2020-11-24T17:15:59-0500

If "R" is a binary relation between the finite sets X and Y, that is "R \u2286 X\u00d7Y", then "R" can be represented by the logical matrix "M_R" whose row and column indices index the elements of "X" and "Y", respectively, such that the entries of "M_R" are defined by:


"{\displaystyle M_{i,j}={\begin{cases}1&(x_{i},y_{j})\in R\\0&(x_{i},y_{j})\not \in R\end{cases}}}"


Since "R\subset A\times A" is reflexive relation, "(x,x)\in R" for all "x\in A". Therefore, "M_{i,i}=1" for all "i\in\{1,...,|A|\}". Consequently, the value of all entries on the main diagonal is 1.